OpenStudy (jolo_yolo):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

Use either elimination or substitution

OpenStudy (anonymous):

Plug and play works too.

OpenStudy (anonymous):

x + 3y = -1 Subtract 3y from both sides x = -1 - 3y Plug in for x 2(-1 - 3y) + 2y = 6 -2 - 6y + 2y = 6 Like terms -2 -4y = 6 Add 2 to both sides -4y = 8 Divide by -4 y = -2

OpenStudy (p0sitr0n):

or you could put it into a matrix and do the Gaussian elimination to obtain the raw echelon form. if you want a fancy way of doing it

OpenStudy (anonymous):

x + 3(-2) = -1 x - 6 = -1 Add 6 to both sides x = 5

OpenStudy (anonymous):

So x = 5 y = -2 (x,y) What is the answer?

OpenStudy (p0sitr0n):

\[\left[\begin{matrix}1 & 3 & |-1\\ 2 & 2 & | \; \; \; \; 6\end{matrix}\right]\]

OpenStudy (jolo_yolo):

(5,-2) Thankss. I guess your explanation made it look easier than I thought it'd be hard

OpenStudy (anonymous):

No problem! :P